OpenStudy (anonymous):

Choose the correct simplification of the expression (5xy^5)^2(y^3)^4. 25x^2y^22 10x^2y^22 25x^3y^14 10x^3y^14

OpenStudy (whpalmer4):

Two properties of exponents needed here: \[(ab)^n = a^nb^n\]\[a^n*a^m = a^{n+m}\] And don't forget that \[a^1 = a\]

OpenStudy (anonymous):

first option

OpenStudy (whpalmer4):

Okay, you also can use \[(a^n)^m = a^{n*m}\]

OpenStudy (whpalmer4):

\[(5xy^5)^2 = (5xy^5)*(5xy^5) = (5*5)*(x*x)*(y^5*y^5) = 25x^2y^{10}\]Repeat the process with the other part and multiply them together to get the final answer.

OpenStudy (whpalmer4):

You could also do it like this: \[(5xy^5)^2 = (5)^2*x^2*(y^5)^2 = 25*x^2*y^{2*5} = 25x^2y^{10}\]

OpenStudy (anonymous):

i dont like u go away. and 25x^2y^10 isnt one of the answer choices so its obviously wrong. btw 1 day and counting and i still havent gotten that explaination. am i fit to have an attitude now ? i think so.

OpenStudy (whpalmer4):

I only explained how to do part of this problem. You need to do the rest. OpenStudy isn't here for you to just get the answers.

OpenStudy (anonymous):

u have already said this. dont repeat yourself like a stupid person. u cause me to repeat MYSELF, which i hate. so like i said before, i want the full explaination. not half. please jus stop "answering" my questions and leave me alone.